Bromination of isobutane is a two-step reaction, as shown below. Using the table of bond dissociation energies, calculate the enthalpy of each step and the enthalpy of the overall reaction.



Bromination of isobutane is a two-step reaction, a



I got the first enthalpy right but the second and the third one wrong. Help!
Answer #1
Concepts and reason

The concept used to solve this problem is based on enthalpy of reaction.

Enthalpy of a reaction is defined as the enthalpy change when stoichiometric quantities of reactants react to form the products in the reaction. The energy required to break a bond is known as bond dissociation energy and the same amount energy released while forming a bond. It may be noted that bond dissociation as well as bond formation take place in chemical reaction.

Fundamentals

Enthalpy of reaction can be calculated by using bond dissociation energy as follow.

ΔH=ΔH(reactants)ΔH(products)\Delta H = \sum \Delta H\left( {{\rm{reactants}}} \right) - \sum \Delta H\left( {{\rm{products}}} \right)

Here, ΔH\Delta H is heat of reaction.

Part 1

The reaction for combustion of acetylene can be written as follow.

(CH3)3CH+Br(CH3)3C+HBr{\left( {{\rm{C}}{{\rm{H}}_3}} \right)_3}{\rm{C}} - {\rm{H}} + {\rm{Br}} \to {\left( {{\rm{C}}{{\rm{H}}_3}} \right)_3}{{\rm{C}}^ \bullet }{\rm{ + H}} - {\rm{Br}}

Enthalpy of reaction for this reaction can be written as follow.

ΔH1=ΔH((CH3)3CH)ΔH(HBr)\Delta {H_1} = \Delta H\left( {{{\left( {{\rm{C}}{{\rm{H}}_3}} \right)}_3}{\rm{C}} - {\rm{H}}} \right) - \Delta H\left( {{\rm{H}} - {\rm{Br}}} \right)

Substitute 400kJmol1400{\rm{ kJ mo}}{{\rm{l}}^{ - 1}} for ΔH((CH3)3CH)\Delta H\left( {{{\left( {{\rm{C}}{{\rm{H}}_3}} \right)}_3}{\rm{C}} - {\rm{H}}} \right) and 366kJmol1{\rm{366 kJ mo}}{{\rm{l}}^{ - 1}} for ΔH(HBr)\Delta H\left( {{\rm{H}} - {\rm{Br}}} \right) .

ΔH1=(400366)kJmol1=+34kJmol1\begin{array}{c}\\\Delta {H_1} = \left( {400 - 366} \right){\rm{ kJ mo}}{{\rm{l}}^{ - 1}}\\\\ = + 34{\rm{ kJ mo}}{{\rm{l}}^{ - 1}}\\\end{array}

Part 2

The reaction for combustion of acetylene can be written as follows:

(CH3)3C+BrBr(CH3)3CBr+Br{\left( {{\rm{C}}{{\rm{H}}_3}} \right)_3}{{\rm{C}}^ \bullet } + {\rm{Br}} - {\rm{Br}} \to {\left( {{\rm{C}}{{\rm{H}}_3}} \right)_3}{\rm{C}} - {\rm{Br}} + {\rm{Br}}

Enthalpy of reaction for this reaction can be written as follows:

ΔH2=ΔH(BrBr)ΔH((CH3)3CBr)\Delta {H_2} = \Delta H\left( {{\rm{Br}} - {\rm{Br}}} \right) - \Delta H\left( {{{\left( {{\rm{C}}{{\rm{H}}_3}} \right)}_3}{\rm{C}} - {\rm{Br}}} \right)

Substitute 292kJmol1{\rm{292 kJ mo}}{{\rm{l}}^{ - 1}} for ΔH((CH3)3CBr)\Delta H\left( {{{\left( {{\rm{C}}{{\rm{H}}_3}} \right)}_3}{\rm{C}} - {\rm{Br}}} \right) and 193kJmol1{\rm{193 kJ mo}}{{\rm{l}}^{ - 1}} for ΔH(BrBr)\Delta H\left( {{\rm{Br}} - {\rm{Br}}} \right) .

ΔH2=(193292)kJmol1=99kJmol1\begin{array}{c}\\\Delta {H_2} = \left( {193 - 292} \right){\rm{ kJ mo}}{{\rm{l}}^{ - 1}}\\\\ = - 99{\rm{ kJ mo}}{{\rm{l}}^{ - 1}}\\\end{array}

Part 3

The reaction for combustion of acetylene can be written as follows:

(CH3)3CH+BrBr(CH3)3CBr+HBr{\left( {{\rm{C}}{{\rm{H}}_3}} \right)_3}{\rm{C}} - {\rm{H}} + {\rm{Br}} - {\rm{Br}} \to {\left( {{\rm{C}}{{\rm{H}}_3}} \right)_3}{\rm{C}} - {\rm{Br}} + {\rm{H}} - {\rm{Br}}

Enthalpy of overall reaction is calculated by adding enthalpy of step 1 and step 2.

ΔH=ΔH1+ΔH2\Delta H = \Delta {H_1} + \Delta {H_2}

Substitute 34kJmol1{\rm{34 kJ mo}}{{\rm{l}}^{ - 1}} for ΔH1\Delta {H_1} and 99kJmol1 - 99{\rm{ kJ mo}}{{\rm{l}}^{ - 1}} for ΔH2\Delta {H_2} .

ΔH=(3499)kJmol1=65kJmol1\begin{array}{c}\\\Delta H = \left( {34 - 99} \right){\rm{ kJ mo}}{{\rm{l}}^{ - 1}}\\\\ = - 65{\rm{ kJ mo}}{{\rm{l}}^{ - 1}}\\\end{array}

Ans: Part 1

Enthalpy of the reaction is +34kJmol1 + 34{\rm{ kJ mo}}{{\rm{l}}^{ - 1}} .